### v1

@@ Abstract
Context:The recent discovery of several ultra high-energy gamma-ray emitters in our Galaxy constitutes a significant advancement towards unveiling its most powerful accelerators and their properties. Nonetheless, in order to unambiguously locate the regions where the highest energy particles are produced and understand the responsible physical mechanisms, detailed spectral and morphological studies are required, especially given that most of the observed sources were found to be significantly extended.
Aims:In these regards, pointing observations with the next-generation Imaging Atmospheric Cherenkov Telescopes, like the Cherenkov Telescope Array (CTA) Observatory and the ASTRI Mini-Array (ASTRI), are expected to provide significant improvements. Here we aim at identifying the most promising sources to target in future observations.
Methods:To this purpose, we performed a comparative analysis of the expected performance of ASTRI and CTA, computing their differential sensitivities towards extended sources, and further explored their capabilities with respect to specific case studies, including follow-ups of existing gamma-ray source catalogs.
Results:We find that almost all of the sources so far detected by LHAASO-WCDA and HGPS will be in the reach of ASTRI and CTA with 300 and 50 hours of exposure, respectively. For the highest energy emitters detected by LHAASO-KM2A, in turn, we provide the list of the most promising objects to be investigated. We further examined specific classes of sources in order to identify potentially detectable gamma-ray emitters, such as passive molecular clouds (i.e. illuminated by the cosmic-ray sea) and pulsars surrounded by a halo of runaway particles.
Conclusions:
Detection prospects of Very and Ultra High-Energy gamma rays from extended sources with ASTRI, CTA, and LHAASO
@@ Conclusions:
In recent years, gamma-ray astronomy has witnessed a significant experimental boost. In the context of the so-called Extensive Atmospheric Shower (EAS) ground-based detectors, the Large High Altitude Air Shower Observatory (LHAASO) started operating in its full configuration in 2021, opening a new observational window into the electromagnetic sky above 100 TeV. Thanks to its large effective area and efficient background rejection, LHAASO achieved in the first year of data acquisition an unprecedented sensitivity, enabling the discovery of few tens of Ultra High-Energy (UHE, i.e. 100 TeV) emitters (Cao et al., 2023). The identification of these source is still uncertain, partly due to the limited angular resolution of its detection technique. In fact, most of the observed sources have a large angular extension and are located in crowded regions of the Galactic Plane, thus increasing the chances for source confusion.
@@
Meanwhile, significant improvements are foreseen regarding Imaging Atmospheric Cherenkov Telescopes (IACTs), with the ongoing construction of the ASTRI Mini-Array (Astrofisica con Specchi a Tecnologia Replicante Italiana, hereafter ASTRI) and the Cherenkov Telescope Array Observatory (CTAO, hereafter CTA). ASTRI is a nine-telescopes array (one telescope is already operative) currently under construction in Tenerife (Spain) (Vercellone et al., 2022a). In its final configuration, which is expected to be completed in the next few years, ASTRI will be sensitive to energies up to 200 TeV with an improved sensitivity compared to the currently operating IACTs (MAGIC, VERITAS and H.E.S.S.) above a few TeV (Lombardi et al., 2022). CTA will in turn consist of two arrays, one located in La Palma (Spain) and the other in Paranal (Chile), thus covering both the Northern and the Southern hemisphere. Once completed, CTA will reach a sensitivity one order of magnitude lower than the currently-operating IACTs, spanning a broad energy interval ranging from a few tens of GeV to a few hundreds of TeV (Cherenkov Telescope Array Consortium et al., 2019).
@@
The advantage of detecting photons with arrays of IACTs is above all their superior angular resolution ( 3 arcmin), peculiar of the stereoscopic technique (Aharonian et al., 2013). This is a key feature for obtaining a clear source identification and to conduct precise morphological investigations, capable of distinguishing the sites of particle acceleration from nearby regions (e.g. molecular clouds illuminated by escaped particles). Furthermore, the larger Field of View (FoV) of the next-generation IACTs (up to in diameter) will allow us to study and resolve large and crowded sources such as supernova remnants, pulsar wind halos, and superbubbles.
@@
Nevertheless, observations suffer from a degraded performance with respect to the ideal on-axis point-like case, due to both off-axis pointing and increased background in extended analyses, which also affects the minimum source flux detectable by IACTs. As shown in Ambrogi et al. (2018) for a previous CTA layout (the so-called 2Q-array), the extent of the worsening depends on the energy range of the observations, and it tends to disappear at the highest energies. Thus, a more accurate evaluation of the detection prospects and required exposure per source is necessary when dealing with extended sources. We here update those predictions by considering the most up-to-date CTA proposed layout, the so-called Alpha configuration111https://www.cta-observatory.org/science/ctao-performance/. In addition, we compute the expected ASTRI and LHAASO sensitivities towards extended sources. Such investigations allow us to clearly define a sample of gamma-ray sources that will be detectable by the next-generation IACTs, which is crucial to outline an optimized observational plan. We start by summarizing the detection techniques of these instruments in Sec. 2, and later describe their specific features in Sec. 3. We then introduce the methods that we developed for computing the instrument sensitivities for extended source observations in Sec. 4, and discuss the obtained results. As an application of the derived minimum detectable fluxes, in Sec. 5 we present cases demanding for tailored IACT observations aimed at obtaining improved morphological and spectral information. Particularly, in Sec. 5.1 we scan existing gamma-ray source catalogs in the Very High-Energy (VHE, i.e. 100 GeV) and UHE bands to define promising candidates for observations, while in the following subsections we discuss perspectives for detecting PeVatrons, passive molecular clouds illuminated by cosmic rays, and TeV halos. We finally draw our conclusions in Sec. 6. Three appendices complement the text: in Appendix A we discuss IACT response functions to off-axis sources, in Appendix B we present IACT differential sensitivities in high-zenith angle observations, and finally in Appendix C we provide a comparison among our results and official point-like source sensitivities from ASTRI, CTA, and LHAASO, also exploring individual sub-arry configurations of CTA.
@@ 2 Detection methods of VHE gamma rays
The Earth atmosphere is opaque to gamma rays, consequently ground-based detection methods rely on the indirect measurements of the secondary particles that compose the cascades generated by photons. Two main detection technique exist for measuring such emission: the IACT technique, based on the observation of the Cherenkov flashes induced by the superluminous shower particles traversing the atmosphere; and the EAS technique, aimed at collecting individual shower particles from ground. The discrimination of photon- from hadron-induced showers is a fundamental part of the measurements. In the following we describe the different detection and background rejection methods implemented by the present and next-generation instruments.
@@ 2.1 IACTs: CTA and ASTRI
Large (1 m) optical reflectors, combined with fast (nanoseconds) cameras sensitive to blue (300-350 nm) light, permit the detection of the Cherenkov emission of the shower ultra-relativistic secondary particles. Because of the signal faintness, IACTs operate only at night, when the Moon contribution is limited, and in good weather conditions, reducing the duty cycle of these detectors ranging in 10-20%. The stereoscopic distribution of the collectors in arrays of multiple telescopes enables a better determination of the arrival direction of the shower, guaranteeing angular resolution of the order of a few arcminutes (Aharonian et al., 2013). Hadron-initiated shower are also detected by IACTs: as they constitute a background for photon-induced cascades, hadronic showers need to be rejected after the data acquisition (Maier & Knapp, 2007). A first gamma/hadron separation is conducted based on the projected shape of the Cherenkov-light cone on the camera (Hillas, 1985): the high transverse momentum of secondary hadrons, in fact, induces a highly scattered distribution of hadronic showers at fixed energy of the primary. Nonetheless, an irreducible background of gamma-ray like hadron will survive the separation cuts. Such a residual hadronic emission is estimated at the level of the analysis, e.g. by evaluating the background in a source-free region or by fitting a background template to the data in the region of interest (Berge et al., 2007). The two methods have been extensively tested for H.E.S.S. data analysis showing equivalent results.
The future of IACTs foresees two new facilities. ASTRI at its completion will be an array of nine telescopes located at the Observatorio del Teide in Tenerife (2817’60.00”N, 16 30’20.99”W) at an altitude of approximately 2370 m. Each of the telescopes will have two mirrors in the Schwarzschild-Couder design, of 4.3 m and 1.8 m diameter respectively, whose size allows the array to best perform in the energy range around 10 TeV. The telescopes are equipped with a wide FoV camera, about in diameter, constituted by silicon photo multipliers (Scuderi et al., 2022; Catalano et al., 2018).
CTA will be an observatory located at two sites: the Northern hemisphere array, so-called CTA-North, designated area is in the Canary Island of La Palma (2845’43.7904”N, 1753’31.218”W), while the Southern array, so-called CTA-South, in the Paranal plateau in Chile (441’0.34”S, 7018’58.84”W). The final CTA layout (the Alpha configuration) has been recently established: it consists of 4 Large-Size Telescopes (LSTs, one of which is already operating) and 9 Medium-Size Telescope (MSTs) for the Northern array, versus 14 MSTs and 37 Small-Size Telescopes (SSTs) for the Southern site. The diameters of the main reflecting surface of the telescopes will be of 23, 11.5, and 4.3 m respectively for the large, medium and small size type, while their FoV will be , , and respectively (Cherenkov Telescope Array Consortium et al., 2019). The use of many SSTs, distributed on a large (3 km) area is expected to improve the detection at high (5-300 TeV) energies, while the employment of the LSTs will allow reaching a better sensitivity at lower (20–150 GeV) energies. Further improvements with respect to the Alpha configuration are expected as a result of the CTA+ program222https://pnrr.inaf.it/progetto-ctaplus/, that has been recently approved and funded for the procurement of two more LSTs and five SSTs specifically for the CTA-South site (Antonelli, 2023).
@@ 2.2 EAS arrays: LHAASO
LHAASO is an EAS observatory built at 4400 m above sea level in the Sichuan province of China. It consists of three type of detectors designed for the study of Cosmic Rays (CRs) and gamma rays across a broad energy range from sub-TeV to beyond the PeV energy scale: the Water Cherenkov Detector Array (WCDA), the Square Kilometer Array (also called KM2A), and the Wide Field-of-view Cherenkov Telescope Array (WFCTA). The WCDA is a 78000 m surface water Cherenkov pond, filled with purified water and equipped with upward facing photomultiplier tubes (PMTs). The KM2A is an array composed of 5195 Electromagnetic Detectors (EDs) and 1188 Muon Detectors (MDs), extending for 1.3 km surface. EDs are plastic scintillators detecting the electromagnetic component of EASs, covered by a lead plate ( cm thick) and a PMT. MDs are in turn cylindric muon tanks, filled with pure water and buried under 2.5 m of soil, as to absorb the electromagnetic component and detect the muonic component of an EAS. The two instruments operate independently, with a core energy range of GeV to TeV for WCDA, and between TeV and PeVs for KM2A. Finally, 12 wide field of view air Cherenkov telescopes compose the WFCTA, complementing the system in CR observations; given that these are not used in gamma-ray related studies, we will not discuss WFCTA further.
The LHAASO operation principle relies on combining information about the collected PMT hit and charge distributions to evaluate shower parameters, as its core location (i.e. the expected position on ground of the primary particle), size (i.e. number of secondary particles), age (related to the amount of traversed matter in the atmosphere) and muon content. From this information, the primary’s arrival direction, energy and nature (hadronic vs electromagnetic) can be determined. Indeed, in the case where the primary is a CR entering the Earth’s atmosphere, the air shower consists of secondary hadrons, muons, leptons, and photons, while in the case of primary photons the cascade is purely electromagnetic, consisting of leptons and photons. In KM2A, the ratio among the measured number of electrons and muons is used to separate CR from gamma-ray induced showers. In WCDA, additional parameters are used for CR rejection, e.g. the shower compactness, which is determined by the measurement of the effective charge beyond a given position from the shower axis (expected to be larger for CR-induced showers).
The observatory was completed in mid 2021: its high sensitivity, particularly at the highest energies probed by KM2A and so far unexplored, was essential to detect the emission of PeV photons from several Galactic accelerators (Cao et al., 2021).
@@ 3 Instrument response functions
A proper comparison among instruments relying on different detection techniques and operating in different energy ranges is only possible through the evaluation of differential sensitivities, establishing the minimum detectable flux in each energy bin. Their computation, presented in the following section, is based on certain parameters, namely the same energy binning, observation time and detection thresholds, as well as on individual Instrument Response Functions (IRFs), specifically angular and energy resolutions, effective area, and expected background rate. We here proceed to the description of the latter quantities, according tho the latest available official releases by each Collaboration.
@@ 3.1 LHAASO IRFs
We adopt the most recent response functions for LHAASO-KM2A, calibrated in the analysis of the Crab Nebula (Dong et al., 2024). We make use specifically of the IRFs relative to observations at zenith angle , to be consistent with those released by ASTRI. Updated responses of the WCDA detector are not yet available, thus we here limit our investigations to KM2A.
We consider the response functions as angular resolution, effective area to gamma rays, and gamma/hadron separation efficiency, all provided at analysis level, i.e. including both cuts for hadron rejection as well as quality cuts selecting well reconstructed events. The IRFs released so far use event selection cuts optimised for point sources as the optimisation for extended sources requires additional care. We expect analyses of extended sources to require more stringent cuts to reduce the background rate, which impacts the effective area. Hence, they require a dedicated investigation by the collaboration itself, that may be available in the next future.
While gamma-ray effective area, angular/energy resolution are provided in the IRFs, the expected background rate can be derived by convolving the all-particle CR flux (taken from Gaisser et al. (2013)) with the effective area to hadronic showers at analysis level. The latter is obtained directly from the available photon effective area, which is first divided by the gamma-ray selection efficiency, and then multiplied for the CR rejection power as provided in Dong et al. (2024). We finally integrate the resulting area with the all-particle CR flux and derive the quantity shown as a yellow line in Fig. 1(b). It is worth noticing that LHAASO-KM2A obtains its high rejection power thanks to the usage of muons for hadron identification: this is particularly important at the highest energies, where the increased abundance of muons facilitates the identification of the nature of primary particles.

### v2

@@ Abstract
Context. The recent discovery of several ultra high-energy gamma-ray emitters in our Galaxy represents a significant advancement towards the characterisation of its most powerful accelerators. Nonetheless, in order to unambiguously locate the regions where the highest energy particles are produced and understand the responsible physical mechanisms, detailed spectral and morphological studies are required, especially given that most of the observed sources were found to be significantly extended.
Aims. In these regards, pointing observations with the next-generation Imaging Atmospheric Cherenkov Telescopes, such as the Cherenkov Telescope Array (CTA) Observatory and the ASTRI Mini-Array (ASTRI), are expected to provide significant improvements. Here we aim to identify the most promising sources to target in future observations.
Methods. For this purpose, we performed a comparative analysis of the expected performance of ASTRI and CTA, computing their differential sensitivities towards extended sources, and further explored their capabilities with respect to specific case studies, including follow-ups of existing gamma-ray source catalogues.
Results. We find that almost all of the sources thus far detected by LHAASO-WCDA and in the H.E.S.S. Galactic Plane Survey will be in the reach of ASTRI and CTA with about 300 and 50 hours of exposure, respectively. For the highest energy emitters detected by LHAASO-KM2A, in turn, we provide a list of the most promising objects that would require further investigation. We additionally examined specific classes of sources in order to identify potentially detectable gamma-ray emitters, such as passive molecular clouds (i.e. illuminated by the cosmic-ray sea) and pulsars surrounded by a halo of runaway particles.
@@
In recent years, gamma-ray astronomy has witnessed a significant experimental boost. Ground-based instruments as HAWC and Tibet-AS have started to survey the sky at very-high energies (VHEs, i.e. 100 GeV 100 TeV) and ultra-high energies (UHEs, i.e. 100 TeV) (Abeysekara et al., 2017c; Amenomori et al., 2019), unveiling the most powerful northern sources. In 2021, the Large High Altitude Air Shower Observatory (LHAASO) started operating in its full configuration: thanks to its unprecedentedly large effective area and efficient background rejection, LHAASO has enabled the discovery of a few tens of UHE emitters (Cao et al., 2024a) in the northern hemisphere, just in the first year of data acquisition, improving on the sensitivity of the currently operating extensive atmospheric shower (EAS) detectors by about one order of magnitude. The planned SWGO and ALPACA detectors (Albert et al., 2019; Sako et al., 2022) aim to survey the southern hemisphere at similar energies, to complement the view of the UHE sky.
@@
Meanwhile, significant developments are foreseen regarding Imaging Atmospheric Cherenkov Telescopes (IACTs), particularly with the ASTRI Mini-Array (Astrofisica con Specchi a Tecnologia Replicante Italiana, hereafter ASTRI) and the Cherenkov Telescope Array Observatory (CTAO, hereafter CTA). ASTRI is a nine-telescope array (one telescope is already operative) currently under construction in Tenerife (Spain) (Vercellone et al., 2022a). In its final configuration, which is expected to be completed in the next few years, ASTRI will be sensitive to energies up to 200 TeV with an improved sensitivity compared to the currently operating IACTs (MAGIC, VERITAS, and H.E.S.S.) above a few TeV (Lombardi et al., 2022). CTA will in turn consist of two arrays, one located in La Palma (Spain) and the other in Paranal (Chile), thus covering both the northern and the southern hemisphere. Once completed, CTA will reach a sensitivity one order of magnitude lower than the currently operating IACTs, spanning a broad energy interval ranging from a few tens of GeV to a few hundred TeV (Cherenkov Telescope Array Consortium et al., 2019).
The identification of the UHE sources observed by HAWC and LHAASO is still uncertain, partly due to the limited angular resolution of their detection technique. In fact, most of the observed sources have a large angular extension and are located in crowded regions of the Galactic Plane, thus increasing the chances for source confusion.
@@
The advantage of detecting photons with arrays of IACTs is above all their superior angular resolution ( 3 arcmin), a specificity of the stereoscopic technique (Aharonian et al., 2013). This is a key feature for obtaining a clear source identification and to conduct precise morphological investigations capable of distinguishing the sites of particle acceleration from nearby regions (e.g. molecular clouds illuminated by escaped particles). Furthermore, the larger Field of View (FoV) of the next-generation IACTs (up to in diameter) will allow us to study and resolve large and crowded sources such as supernova remnants, pulsar wind halos, and superbubbles.
@@
Nevertheless, IACT observations suffer from a degraded performance with respect to the ideal on-axis point-like case, due to both off-axis pointing and increased background in extended analyses, which also affects their minimum detectable source flux. As shown in Ambrogi et al. (2018) for a previous CTA layout (the so-called 2Q-array), the extent of the worsening depends on energy, and it tends to disappear at the highest energies. Thus, a more accurate evaluation of the detection prospects and required exposure per source is necessary when dealing with extended sources. For this work, we revised those predictions by considering the most up-to-date CTA proposed layout, the so-called Alpha configuration111https://www.cta-observatory.org/science/ctao-performance/, and we provide a new computation of the CTA differential sensitivity for extended sources. In addition, we compute the expected ASTRI and LHAASO differential sensitivities towards extended sources. Such investigations allow us to clearly define a sample of gamma-ray sources that are foreseen to be detectable by the next-generation IACTs, and therefore it will be crucial to outline an optimised observational plan. We start by summarising the detection techniques of these instruments in Sec. 2, and later describe their specific features in Sec. 3. We then introduce the methods that we developed for computing instrumental sensitivities towards extended source observations in Sec. 4, and discuss the obtained results. As an application of the derived minimum detectable fluxes, in Sec. 5 we explain how we scanned existing gamma-ray source catalogues in the VHE and UHE bands to define promising candidates for observations, demanding tailored IACT observations aimed at obtaining improved morphological and spectral information. Particularly, we discuss perspectives for detecting PeVatrons, passive molecular clouds illuminated by cosmic rays, and TeV halos. We then summarise the most promising targets in Sec. 6, and we finally draw our conclusions in Sec. 7. Four appendices complement the text: in Appendix A, we show the integral sensitivities of the considered instruments; in Appendix B we discuss IACT response functions to off-axis sources, in Appendix C we present IACT differential sensitivities in high-zenith angle observations; and finally in Appendix D we provide a comparison among our results and official point-like source sensitivities from ASTRI, CTA, and LHAASO, also exploring individual sub-array configurations of CTA.
@@ 2 Detection methods of VHE gamma rays
The Earth atmosphere is opaque to gamma rays, consequently ground-based detection methods rely on the indirect measurements of the secondary particles that compose the cascades generated by primary photons. Two main detection techniques exist for measuring such emission: the IACT technique, based on the observation at night of the Cherenkov flashes induced by the superluminous shower particles traversing the atmosphere; and the EAS technique, aimed at collecting individual shower particles from ground. The discrimination of photon- from hadron-induced showers is a fundamental part of the measurements. In the following we briefly describe the different detection and background rejection methods implemented by the present and next-generation instruments, referring to the dedicated works from each Collaboration for further details.
@@ 2.1 IACTs: CTA and ASTRI
The future of IACTs foresees two new facilities. ASTRI at its completion will be an array of nine telescopes located at the Observatorio del Teide in Tenerife (28∘17’60.00”N, 16∘ 30’20.99”W) at an altitude of approximately 2370 m. Each of the telescopes will have two mirrors in the Schwarzschild-Couder design, of 4.3 m and 1.8 m diameter respectively, whose size allows the array to best perform in the energy range around 10 TeV. The telescopes are equipped with a wide FoV camera, about in diameter, constituted by silicon photo multipliers (Scuderi et al., 2022; Catalano et al., 2018).
CTA will be an observatory located at two sites: the northern hemisphere array, so-called CTA-North, whose designated area is in the Canary Island of La Palma (28∘45’43.7904”N, 17∘53’31.218”W), while the southern array, so-called CTA-South, will be in the Paranal plateau in Chile (4∘41’0.34”S, 70∘18’58.84”W). The final CTA layout (the Alpha configuration) has been recently established: it consists of 4 Large-Size Telescopes (LSTs, one of which is already operating) and 9 Medium-Size Telescope (MSTs) for the northern array, versus 14 MSTs and 37 Small-Size Telescopes (SSTs) for the southern site. The diameters of the main reflecting surface of the telescopes will be of 23, 11.5, and 4.3 m respectively for the large, medium and small size type, while their FoV will be , , and respectively (Cherenkov Telescope Array Consortium et al., 2019). The use of many SSTs, distributed on a large (3 km2) area, is expected to improve the detection at high (5-300 TeV) energies, while the employment of the LSTs will allow a better sensitivity to be reached at lower (20–150 GeV) energies. Further improvements with respect to the Alpha configuration are expected as a result of the CTA+ program222https://pnrr.inaf.it/progetto-ctaplus/, that has been recently approved and funded for the procurement of two more LSTs and five SSTs specifically for the CTA-South site (Antonelli, 2023).
@@ 2.2 EAS arrays: LHAASO
LHAASO (LHAASO collaboration, 2021; Aharonian et al., 2021a) is an EAS observatory built at 4400 m above sea level in the Sichuan province of China. It consists of three type of detectors designed for the study of Cosmic Rays (CRs) and gamma rays across a broad energy range from sub-TeV to beyond the PeV energy scale: the Water Cherenkov Detector Array (WCDA), the Square Kilometer Array (also called KM2A), and the Wide Field-of-view Cherenkov Telescope Array (WFCTA). The WCDA is a 78000 m2 surface water Cherenkov pond, filled with purified water and equipped with upward facing photomultiplier tubes (PMTs). The KM2A is an array composed of 5195 Electromagnetic Detectors (EDs) and 1188 Muon Detectors (MDs), extending for 1.3 km2 surface. EDs are plastic scintillators detecting the electromagnetic component of EASs, covered by a lead plate ( cm thick) and a PMT. MDs are in turn cylindric muon tanks, filled with pure water and buried under 2.5 m of soil, as to absorb the electromagnetic component and detect the muonic component of an EAS. These two instruments operate independently, with a core energy range of GeV to TeV for WCDA, and between TeV and PeVs for KM2A. Finally, 12 wide field of view air Cherenkov telescopes compose the WFCTA, complementing the system in CR observations; given that these are not used in gamma-ray related studies, we do not discuss WFCTA further. The observatory high sensitivity, particularly at the highest energies probed by KM2A and so far unexplored, was essential to detect the emission of PeV photons from several Galactic accelerators (Cao et al., 2021).
@@ 3 Instrument response functions
Comparing instruments relying on different detection techniques and operating in different energy ranges is only possible through the evaluation of differential sensitivities, which establish, independently for every instrument, the minimum detectable flux per energy bin (Aharonian, 1991). The differential sensitivity computation, presented in the following section, is based on certain parameters, namely the same energy binning, observation time and detection thresholds, as well as on individual Instrument Response Functions (IRFs), specifically angular and energy resolutions, effective area, and expected background rate. We here proceed to the description of the latter quantities, according tho the latest available official releases by each Collaboration.
@@ 3.1 LHAASO IRFs
We adopt the most recent response functions for LHAASO-KM2A, calibrated in the analysis of the Crab Nebula (Cao et al., 2024b). We make use specifically of the IRFs relative to observations at zenith angle , to be consistent with those released by all of the experiments here investigated. Up-to-date performance of the WCDA detector is not yet available, thus we here limit our discussion to KM2A.
We consider the response functions as angular resolution, effective area to gamma rays, and gamma/hadron separation efficiency, all provided at analysis level, that is, including both cuts for hadron rejection as well as quality cuts selecting well reconstructed events. The IRFs released so far use event selection cuts optimised for point sources as the optimisation for extended sources requires additional care. We expect analyses of extended sources to require more stringent cuts to reduce the background rate, which impacts the effective area. Hence, they require a dedicated investigation by the collaboration itself, that may be available in the next future.
While gamma-ray effective area, angular/energy resolution are provided in the IRFs, the expected background rate can be derived by convolving the all-particle CR flux (taken from Gaisser et al. (2013)) with the effective area to hadronic showers at analysis level. The latter is obtained directly from the available photon effective area, which is first divided by the gamma-ray selection efficiency, and then multiplied for the CR rejection power as provided in Cao et al. (2024b). We finally integrate the resulting area with the all-particle CR flux and derive the quantity shown as a yellow line in Fig. 1(b). It is worth noticing that LHAASO-KM2A obtains its high rejection power thanks to the usage of muons for hadron identification: this is particularly important at the highest energies, where the increased abundance of muons facilitates the identification of the nature of primary particles.